OpenStudy (anonymous):

If five less than three-fourths of an integer is the same as five more than one-eighth of the same integer, what is the integer?

OpenStudy (hoblos):

3/4 n -5 = 1/8 n +5 => 6/8 n - 1/8 n = 10 => 5/8 n = 10 => n= 16